If a selected home takes under 30 days or over 90 days to sell, what is the probability its initial asking price is under $150,000 or over $250,000 Question 4 Question 5 What is the probability a selected home takes at most days to sell or its initial asking price is under $200,000? What is the probability a selected home's initial asking price is between $150,000 and $250,000? Question Question 7A What is the probability a selected home's initial asking price is at least $150.000 and it takes between 30 and 90 days to sell?
KEY Price1 Under $150,000 Days1 Under 30 - Note: Pivot table MUST start in Price2 $150,000 $199,999 Days2 30-90 Price3 $200,000 $250,000 Days3 Over 90 cell D8 of this sheet. Price4 Over $250,000 Count of Days Price 4 Grand Total Days 1 2 3 73 1 53 71 66 263 2 56 73 74 54 257 3 69 73 64 74 280 Grand Total 178 219 209 194 800
Answer #1

4)
P( price 1 OR 4 | day 1 OR 3)
P( price 1 OR 4 | day 1) + P( price 1 OR 4 | day 3)
=(66+53)/263+(69+74)/280
0.9632

5)
P(day 1&2&3 OR price 1&2)
P(day 1&2&3 with price 1) + P(day 1&2&3 with price2 )
=(178+219)/800
0.49625

6)
P(PRICE 2 or 3)
=(219+209)/800
0.535

7)
P(price>=1 AND day1&2&3)
=(178+219+209+194)/800
1